A thick spherical shell has inner radius a and outer b, the material has resistivity r. when a potential difference is applied between the inner & outer surfaces, its resistance is?

Grade:11

1 Answers

Khimraj
6 years ago
We assume a shell element of thickness dp and of radius p.
then resistance is given by
dR = rdp/4πp2
R = (r/4π)\int_{a}^{b}dp/p2 = (r/4π)(1/a – 1/b).
Hope it clears. If you like anser then please approve it.